OpenStudy (anonymous):

What is 6 • 9 • (–5)?

OpenStudy (anonymous):

Use Order Of Operations. PEMDAS

OpenStudy (anonymous):

-270

OpenStudy (anonymous):

Do you have a calculator on your computer???

OpenStudy (anonymous):

Whats with all the kindergarden math lately ?, have you tried a calculator, to check your answer ?

OpenStudy (anonymous):

all you have to do is just multiply the numbers and a positive times a negative = a negative!

OpenStudy (anonymous):

Thats what I would put...I'm not sure. I tried it on my computer and thats what it said..

OpenStudy (anonymous):

Thanks I was just checking with you guys because this is a big test im on

OpenStudy (anonymous):

It's just multiplication, if you find it hard remembering all the tables, just do it stepwise and simplify the expression

OpenStudy (anonymous):

I do remember its just this is a semester test is 50 percent of my grade you see

OpenStudy (anonymous):

for example it's faster for me to do it like this: 6*10-6=54 now 54*-5=-250-20=-270